### v2

@@ 3.1 The Fokker-Planck equation
From now on, we fix $m\in\mathbb{N}$ (to be specified later) and a constant $\alpha>0$. Since we will need to work with diffusion coefficients which are never trivial, for every $(a,b)\in\mathcal{H}^{+}_{m}$ we will rather replace $a$ with the mapping
$a+\alpha I:[0,T]\times\mathbb{R}^{n}\to\textnormal{Sym}_{++}(n)\quad\textnormal{with}\quad a+\alpha I=\sqrt{a+\alpha I}\sqrt{a+\alpha I}.$
@@
$\textnormal{SDE}_{x}\quad\begin{cases}\mathrm{d}X_{x}(t)=b(t,X_{x}(t))\;\mathrm{d}t+\sqrt{(a+\alpha I)(t,X_{x}(t))}\;\mathrm{d}W_{t},\\[5.0pt] \mathbb{P}\big{(}X_{x}(0)=x\big{)}=1,\end{cases}$
holds in $(\Omega,\mathcal{G},\mathcal{F},\mathbb{P})$ for $\mu_{0}$-almost every $x\in\mathbb{R}^{n}$. We then say that $X$ solves or is solution to SDE (with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$). The solution $X$ to SDE is unique in the following sense: if a measurable mapping $Y:\mathbb{R}^{n}\times\Omega\to S$ solves SDE, then it holds that $X(x,\cdot)=Y(x,\cdot)$ a.s., for $\mu_{0}$-almost every $x\in\mathbb{R}^{n}$.
For the well-posedness of Theorem-Definition 3.1, see Section 8. Solutions to SDE share a close relationship with the solutions to the Fokker-Planck equation, which we introduce next. For every $(a,b)\in\mathcal{H}^{+}_{m}$, we denote the Kolmogorov generator by
$\mathcal{L}^{a,b}_{t}\varphi(y)\triangleq\frac{1}{2}\sum^{n}_{i,j=1}(a+\alpha I)_{ij}(t,y)\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}(y)+\sum^{n}_{i=1}b_{i}(t,y)\frac{\partial\varphi}{\partial y_{i}}(y),\quad\varphi\in C^{2}(\mathbb{R}^{n},\mathbb{R}).$
Let $X:\mathbb{R}^{n}\times\Omega\to S$ be solution of SDE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$. Thanks to specific regularity properties of $X$ (we refer to Section 8 for these latter), we may define the curve $\mu:[0,T]\to\mathcal{P}(\mathbb{R}^{n})$ of probability measures
@@
We then say that $\mu$ solves or is solution to FPE (with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$). If $X$ denotes the solution to SDE, the following representation formula holds:
@@ 3.2 Absolutely continuous solutions to the Fokker-Planck equation
Thanks to the regularity of the coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ which is offered through Theorem 2.1, the solution $\mu$ to FPE is absolutely continuous, i.e., it takes the form
@@
Let $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ and $\bar{p}\in L^{2}(\mathbb{R}^{n},\mathbb{R})$. A (regular enough) function $p:[0,T]\times\mathbb{R}^{n}\to\mathbb{R}$ is said to solve or be solution to the non-homogeneous Fokker-Planck Equation (with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$), if
@@
For every $(a,b)\in\mathcal{H}^{+}_{m}$, $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$, and $\bar{p}\in L^{2}(\mathbb{R}^{n},\mathbb{R})$, there exists a unique mapping $p\in C(0,T;L^{2}(\mathbb{R}^{n},\mathbb{R}))\cap H^{0,1}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ which solves FPEf with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$. In addition, the following estimate holds:
$\|p(t,\cdot)\|^{2}_{L^{2}}+\int^{t}_{0}$ $\|p(t,\cdot)\|^{2}_{H^{1}}\;\mathrm{d}t\leq$
(9) $\leq C\big{(}\|(a,b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{L^{2}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right),\quad t\in[0,T],$
where $C\big{(}\|(a,b)\|_{\mathcal{H}_{m}}\big{)}>0$ is a constant which continuously depends on $\|(a,b)\|_{\mathcal{H}^{+}_{m}}$.
@@
and the following Strong Fokker-Planck Equations (with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$):
@@
Combining Theorem 3.2 with the representation formula (8) allows us to introduce criteria to establish satisfactory guarantees for our learning approach in the context of observation and regulation of stochastic differential equations (a more exhaustive presentation is provided in Section 8). More specifically, let $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ be an observation/regulation integral metric and $(a,b)\in\mathcal{H}^{+}_{m}$. By denoting $X^{a,b}$ and $p_{a,b}$ respectively the solution to SDE and to SFPE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, we are interested in studying the accuracy with which the following observation/regulation metric is approximated by our learning approach:
@@
For every $(a_{1},b_{1}),(a_{2},b_{2})\in\mathcal{H}^{+}_{m}$, it holds that
@@ 4 Useful Results from RKHS Theory
Let $\ell\in\mathbb{N}$, $D\subseteq\mathbb{R}^{s}$ be an open and bounded domain, and $u\in H^{r}(D,\mathbb{R})$ be a function such that $u|_{\widehat{X}}=0$, where $\widehat{X}_{\ell}\triangleq\{x_{1},\dots,x_{\ell}\}\subseteq\mathbb{R}^{s}$ is a finite set of $\ell$ given points. For the fill distance, which is defined by
@@ 5 Learning Stochastic Differential Equations
    Let $m\in\mathbb{N}$ with $m\geq 1$, $\alpha>0$, and $p_{0}\in H^{2m+1}(\mathbb{R}^{n},\mathbb{R})$. There exist $(a_{*},b_{*})\in\mathcal{H}^{+}_{m}$ and a (unique) solution $X$ to SDE with coefficients $(a_{*},b_{*})\in\mathcal{H}^{+}_{m}$.
@@
    Second, we approximate the unknown coefficients $(a_{*},b_{*})\in\mathcal{H}^{+}_{m}$ through additional coefficients $(\widehat{a},\widehat{b})\in\mathcal{H}^{+}_{m}$ which “best” match SFPE when evaluated at $\widehat{p}$, through an appropriately well-posed infinite-dimensional learning problem. We then combine appropriate estimates with (3.2) to show the solution to SFPE with coefficients $(\widehat{a},\widehat{b})$ well-approximates the unknown $p$.
@@ 5.1 Approximating solutions to SDE via RKHS-based models
The main benefit offered by replacing the unknown density $p$ with the model density $\widehat{p}$ consists of the fact that integrals of this latter mapping and of its derivatives can be actually (easily) computed, enabling to correctly instantiate our first learning problem for $(a_{*},b_{*})\in\mathcal{H}^{+}_{m}$, which is our next step.
@@ 5.2 The infinite-dimensional learning problem and its fidelity
In this section, our goal consists of instantiating and analyzing the learning problem to identify the coefficients $(a_{*},b_{*})\in\mathcal{H}^{+}_{m}$. This problem computes coefficients which “best” matches SFPE when evaluated at the model density $\widehat{p}$. Below, we seek those coefficients in the whole set $\mathcal{H}^{+}_{m}$, deferring to a later section the problem of learning $(a_{*},b_{*})\in\mathcal{H}^{+}_{m}$ through finite-dimensional models. In particular, the main result contained herein consists of appropriate learning error estimates which pave the way to obtaining learning error estimates for the problem of learning $(a_{*},b_{*})\in\mathcal{H}^{+}_{m}$ through finite-dimensional models. However, for pedagogical purposes, we present complete learning rates for the infinite-dimensional learning problem as well.
@@
    There exists $R_{*}>0$ such that the coefficients $(a_{*},b_{*})\in\mathcal{H}^{+}_{m}$ satisfy:
    $\textnormal{supp}\big{(}a_{*}(t,\cdot),b_{*}(t,\cdot)\big{)}\subseteq\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)},\quad\textnormal{for every}\ t\in[0,T].$
@@
Assumption $(B)$ plays a key role in computing learning rates which, on the one hand, leverage classical RKHS approximation theory, and which, on the other hand, depend explicitly on the parameters defining the learning problem. In particular, the latter property can not be generally obtained by just (smoothly) restricting to balls quantities which are defined on unbounded domains. That said, our result can be shown to hold under less restricting assumptions, such as assuming the coefficients $(a_{*},b_{*})\in\mathcal{H}^{+}_{m}$ decay to zero at infinity under specific rates [48], although we opted for Assumption $(B)$ to avoid excessively tedious computations, in turn fostering a smooth exposition. In addition, it is worth mentioning Assumption $(B)$ is often naturally verified in many applications ranging from biology to robotics, where the state space is a bounded domain.
Thanks to Assumption $(B)$, we may restrict ourselves to the closed subspace:
$\mathcal{H}^{+}_{m,R_{*}}\triangleq\Big{\{}(a,b)\in\mathcal{H}^{+}_{m}:\ \textnormal{supp}\big{(}a(t,\cdot),b(t,\cdot)\big{)}\subseteq\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)},\ \textnormal{for every}\ t\in[0,T]\Big{\}}\subseteq\mathcal{H}^{+}_{m},$
We are now ready to define our infinite-dimensional problem for learning the coefficients $(a_{*},b_{*})\in\mathcal{H}^{+}_{m,R_{*}}$, which writes as follows:
@@
$\textnormal{LP}\quad\underset{(a,b)\in\mathcal{H}^{+}_{m,R_{*}}}{\min}L_{\lambda}(a,b)\triangleq\int^{T}_{0}\left\|\frac{\partial\widehat{p}}{\partial t}(t,\cdot)-(\mathcal{L}^{a,b}_{t})^{*}\widehat{p}(t,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}t+\lambda\|(a,b)\|^{2}_{\mathcal{H}_{m}}\ .$
@@
For every $\lambda>0$, problem LP is well-posed and has a unique solution, which is denoted by $(\widehat{a},\widehat{b})\in\mathcal{H}^{+}_{m,R_{*}}$.
@@
(18) $(a,b)\in\mathcal{H}_{m}\mapsto\int^{T}_{0}\left\|\frac{\partial\widehat{p}}{\partial t}(t,\cdot)-(\mathcal{L}^{a,b}_{t})^{*}\widehat{p}(t,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}t+\lambda\|(a,b)\|^{2}_{\mathcal{H}_{m}}$
is strictly convex when restricted to the closed convex subset $\mathcal{H}^{+}_{m,R_{*}}\subseteq\mathcal{H}_{m}$, we just need to prove the existence of a solution to LP. For this, if $(a_{k},b_{k})_{k\in\mathbb{N}}\in\mathcal{H}^{+}_{m,R_{*}}$ is any minimizing sequence for LP, there must exist some constant $C>0$ such that $\|(a_{k},b_{k})\|_{\mathcal{H}^{+}_{m}}\leq C$ for every $k\in\mathbb{N}$, and therefore, since $\mathcal{H}^{+}_{m,R_{*}}$ is in particular a closed and convex subset of the Hilbert space $\mathcal{H}_{m}$, up to extracting a subsequence there exists $(\widehat{a},\widehat{b})\in\mathcal{H}^{+}_{m,R_{*}}$ such that $(a_{k},b_{k})_{k\in\mathbb{N}}$ converges to $(\widehat{a},\widehat{b})$ for the weak topology of $\mathcal{H}_{m}$. Now, thanks to Theorem 2.1, it is clear the mapping (18) is in addition continuous for the strong topology of $\mathcal{H}_{m}$. Thus, we infer that the mapping (18) is in particular weakly lower semi-continuous, and therefore we finally obtain that
$\int^{T}_{0}\left\|\frac{\partial\widehat{p}}{\partial t}(t,\cdot)-(\mathcal{L}^{\widehat{a},\widehat{b}}_{t})^{*}\widehat{p}(t,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}t+\lambda\|(\widehat{a},\widehat{b})\|^{2}_{\mathcal{H}_{m}}\leq$
$\leq\underset{k\to\infty}{\liminf}\ \int^{T}_{0}\left\|\frac{\partial\widehat{p}}{\partial t}(t,\cdot)-(\mathcal{L}^{a_{k},b_{k}}_{t})^{*}\widehat{p}(t,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}t+\lambda\|(a_{k},b_{k})\|^{2}_{\mathcal{H}_{m}}$
$=\underset{(a,b)\in\mathcal{H}^{+}_{m,R_{*}}}{\min}\quad\int^{T}_{0}\left\|\frac{\partial\widehat{p}}{\partial t}(t,\cdot)-(\mathcal{L}^{a,b}_{t})^{*}\widehat{p}(t,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}t+\lambda\|(a,b)\|^{2}_{\mathcal{H}_{m}},$
@@
where the constant $C(a_{*},b_{*})>0$ depends on $a_{*}$ and $b_{*}$ uniquely.
@@
$L_{\lambda}(\widehat{a},\widehat{b})-$ $\lambda\|(a_{*},b_{*})\|^{2}_{\mathcal{H}_{m}}\leq$
@@
$\leq C(a_{*},b_{*})L(\widehat{p})\leq C(a_{*},b_{*}){\color[rgb]{0,0,0}\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2}},$
@@
We better formalize this metric as follows. For every $(a,b)\in\mathcal{H}^{+}_{m}$, let $p_{a,b}\in H^{m+1,2(m+1)}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ denote the unique solution to SFPE with coefficients $(\widehat{a},\widehat{b})\in\mathcal{H}^{+}_{m}$. Note that the existence and uniqueness of the regular mapping $p_{a,b}\in H^{m+1,2(m+1)}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ as non-negative solution to SFPE, of unitary mass and with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ is immediate consequence of Theorem 3.2. We define the following metric to test the accuracy of our method:
$E(a,b)\triangleq\|p_{a,b}-p\|^{2}_{L^{2}}=\int^{T}_{0}\|p_{a,b}(t,\cdot)-p(t,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}t,\quad(a,b)\in\mathcal{H}^{+}_{m}.$

### v3

@@ 3.1 The Fokker-Planck equation
From now on, we fix $m\in\mathbb{N}$ and a constant $\alpha>0$ (to be specified later). Since we will need to work with diffusion coefficients that are never trivial, we replace $\mathcal{H}^{+}_{m}$ with
${\color[rgb]{0,0,0}{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}\triangleq(\alpha I,0)+\Big{\{}(a,b)\in\mathcal{H}^{+}_{m}:\ a(t,x)\succcurlyeq 0,\ (t,x)\in[0,T]\times\mathbb{R}^{n}\Big{\}}.}$
@@
$\textnormal{SDE}_{x}\quad\begin{cases}\mathrm{d}X_{x}(t)=b(t,X_{x}(t))\;\mathrm{d}t+{\color[rgb]{0,0,0}\sqrt{a(t,X_{x}(t))}}\;\mathrm{d}W_{t},\\[5.0pt] \mathbb{P}\big{(}X_{x}(0)=x\big{)}=1,\end{cases}$
holds in $(\Omega,\mathcal{G},\mathcal{F},\mathbb{P})$ for $\mu_{0}$-almost every $x\in\mathbb{R}^{n}$. We then say that $X$ solves or is solution to SDE (with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$). The solution $X$ to SDE is unique in the following sense: if a measurable mapping $Y:\mathbb{R}^{n}\times\Omega\to S$ solves SDE, then it holds that $X(x,\cdot)=Y(x,\cdot)$ a.s., for $\mu_{0}$-almost every $x\in\mathbb{R}^{n}$.
For the well-posedness of Theorem-Definition 3.1, see Section 8. Solutions to SDE share a close relationship with the solutions to the Fokker-Planck equation, which we introduce next. For every $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, we denote the Kolmogorov generator by
$\mathcal{L}^{a,b}_{t}\varphi(y)\triangleq\frac{1}{2}\sum^{n}_{i,j=1}{\color[rgb]{0,0,0}a}_{ij}(t,y)\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}(y)+\sum^{n}_{i=1}b_{i}(t,y)\frac{\partial\varphi}{\partial y_{i}}(y),\quad\varphi\in C^{2}(\mathbb{R}^{n},\mathbb{R}).$
Let $X:\mathbb{R}^{n}\times\Omega\to S$ be solution of SDE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. Thanks to specific regularity properties of $X$ (we refer to Section 8 for these latter), we may define the curve $\mu:[0,T]\to\mathcal{P}(\mathbb{R}^{n})$ of probability measures
@@
We then say that $\mu$ solves or is solution to FPE (with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$). If $X$ denotes the solution to SDE, the following representation formula holds:
@@ 3.2 Absolutely continuous solutions to the Fokker-Planck equation
Thanks to the regularity of the coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ which is offered through Theorem 2.1, the solution $\mu$ to FPE is absolutely continuous, i.e., it takes the form
@@
Let $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ and $\bar{p}\in L^{2}(\mathbb{R}^{n},\mathbb{R})$. A (regular enough) function $p:[0,T]\times\mathbb{R}^{n}\to\mathbb{R}$ is said to solve or be solution to the non-homogeneous Fokker-Planck Equation (with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$), if
@@
For any $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$, and $\bar{p}\in L^{2}(\mathbb{R}^{n},\mathbb{R})$, there exists a unique mapping $p\in C(0,T;L^{2}(\mathbb{R}^{n},\mathbb{R}))\cap H^{0,1}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ which solves FPEf with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. In addition, the following estimate holds:
$\|p(t,\cdot)\|^{2}_{L^{2}}+\int^{t}_{0}$ $\|p(s,\cdot)\|^{2}_{H^{1}}\;\mathrm{d}s\leq$
(9) $\leq C\big{(}{\color[rgb]{0,0,0}\alpha},\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{L^{2}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right),\quad t\in[0,T],$
with $C\big{(}{\color[rgb]{0,0,0}\alpha},\|(a-\alpha I,b)\|_{\mathcal{H}_{m}}\big{)}>0$ continuously depending on ${\color[rgb]{0,0,0}\alpha}$ and $\|(a-\alpha I,b)\|_{\mathcal{H}_{m}}$.
@@
and the following Strong Fokker-Planck Equations (with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$):
@@
Combining Theorem 3.2 with the representation formula (8) allows us to introduce criteria to establish satisfactory guarantees for our learning approach in the context of observation and regulation of stochastic differential equations (a more exhaustive presentation is provided in Section 8). More specifically, let $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ be an observation/regulation integral metric and $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. By denoting $X^{a,b}$ and $p_{a,b}$ respectively the solution to SDE and to SFPE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, we are interested in studying the accuracy with which the following observation/regulation metric is approximated by our learning approach:
@@
For every $(a_{1},b_{1}),(a_{2},b_{2})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, it holds that
@@ 4 Useful Results from RKHS Theory
Let $\ell\in\mathbb{N}$, $D\subseteq\mathbb{R}^{s}$ be a bounded domain, and $u\in H^{r}(D,\mathbb{R})$ be a function such that $u|_{\widehat{X}}=0$, where $\widehat{X}_{\ell}\triangleq\{x_{1},\dots,x_{\ell}\}\subseteq\mathbb{R}^{s}$ is a finite set of $\ell$ given points. For the fill distance, which is defined by
@@ 5 Learning Stochastic Differential Equations
    Let $m\in\mathbb{N}$ with $m\geq 1$, $\alpha>0$, and $p_{0}\in H^{2m+1}(\mathbb{R}^{n},\mathbb{R})$ (these latter will be considered as hyper-parameters in this work). There exist $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ and a (unique) solution $X$ to SDE with coefficients $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$.
@@
    Second, we approximate the unknown coefficients $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ through additional coefficients $(\widehat{a},\widehat{b})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ which “best” match SFPE when evaluated at $\widehat{p}$, through an appropriately well-posed infinite-dimensional learning problem. We then combine appropriate estimates with (3.2) to show the solution to SFPE with coefficients $(\widehat{a},\widehat{b})$ well-approximates the unknown $p$.
@@ 5.1 Approximating solutions to SDE via RKHS-based models
The main benefit offered by replacing the unknown density $p$ with the model density $\widehat{p}$ consists of the fact that integrals of this latter mapping and of its derivatives can be actually (easily) computed, enabling to correctly instantiate our first learning problem for $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, which is our next step.
@@ 5.2 The infinite-dimensional learning problem and its fidelity
In this section, our goal consists of instantiating and analyzing the learning problem to identify the coefficients $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. This problem computes coefficients which “best” matches SFPE when evaluated at the model density $\widehat{p}$. Below, we seek those coefficients in the whole set ${\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, deferring to a later section the problem of learning $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ through finite-dimensional models. In particular, the main result contained herein consists of appropriate learning error estimates which pave the way to obtaining learning error estimates for the problem of learning $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ through finite-dimensional models. However, for pedagogical purposes, we present complete learning rates for the infinite-dimensional learning problem as well.
@@
    There exists $R_{*}>0$ such that the coefficients $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ satisfy:
    $\textnormal{supp}\big{(}{\color[rgb]{0,0,0}a_{*}(t,\cdot)-\alpha I},b_{*}(t,\cdot)\big{)}\subseteq\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)},\quad\textnormal{for every}\ t\in[0,T].$
@@
Assumption $(B)$ plays a key role in computing learning rates which, on the one hand, leverage classical RKHS approximation theory, and which, on the other hand, depend explicitly on the parameters defining the learning problem. In particular, the latter property can not be generally obtained by just (smoothly) restricting to balls quantities which are defined on unbounded domains. That said, our result can be shown to hold under less restricting assumptions, such as assuming the coefficients $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ decay to zero (up to adding $(\alpha I,0)$ of course) at infinity under specific rates [48], although we opted for Assumption $(B)$ to avoid excessively tedious computations, in turn fostering a smooth exposition. In addition, it is worth mentioning Assumption $(B)$ is often naturally verified in many applications ranging from biology to robotics, where the state space is a bounded domain.
Thanks to Assumption $(B)$, we may restrict ourselves to the subset:
${\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}\triangleq\Big{\{}(a,\,$ $b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}:$
$\textnormal{supp}\big{(}{\color[rgb]{0,0,0}a(t,\cdot)-\alpha I},b(t,\cdot)\big{)}\subseteq\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)},\ \textnormal{for every}\ t\in[0,T]\Big{\}}\subseteq{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}},$
We are now ready to define our infinite-dimensional problem for learning the coefficients $(a_{*},b_{*})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}$, which writes as follows:
@@
$\textnormal{LP}\quad\underset{(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}}{\min}L_{\lambda}(a,b)\triangleq\int^{T}_{0}\left\|\frac{\partial\widehat{p}}{\partial t}(t,\cdot)-(\mathcal{L}^{a,b}_{t})^{*}\widehat{p}(t,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}t+\lambda\|({\color[rgb]{0,0,0}a-\alpha I},b)\|^{2}_{\mathcal{H}_{m}}\ .$
@@
For every $\lambda>0$, problem LP is well-posed and has a unique solution, which is denoted by $(\widehat{a},\widehat{b})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}$.
@@
(18) $(a,b)\in\mathcal{H}_{m}\mapsto\int^{T}_{0}\left\|\frac{\partial\widehat{p}}{\partial t}(t,\cdot)-({\color[rgb]{0,0,0}\mathcal{L}^{a+\alpha I,b}_{t}})^{*}\widehat{p}(t,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}t+\lambda\|(a,b)\|^{2}_{\mathcal{H}_{m}}$
is strictly convex on the closed convex subset
${\color[rgb]{0,0,0}\mathcal{S}\triangleq\Big{\{}(a,b)\in\mathcal{H}^{+}_{m}:}$ ${\color[rgb]{0,0,0}a(t,x)\succcurlyeq 0,}$
${\color[rgb]{0,0,0}\textnormal{supp}\big{(}{\color[rgb]{0,0,0}a(t,\cdot)},b(t,\cdot)\big{)}\subseteq\overline{B^{\mathbb{R}^{n}}_{R_{*}}(0)},\quad(t,x)\in[0,T]\times\mathbb{R}^{n}\Big{\}}\subseteq\mathcal{H}_{m}},$
we just need to prove the existence of a solution to LP. For this, if $(a_{k}+\alpha I,b_{k})_{k\in\mathbb{N}}$ is any minimizing sequence for LP, there must exist some constant $C>0$ such that $\|(a_{k},b_{k})\|_{\mathcal{H}_{m}}\leq C$ for every $k\in\mathbb{N}$, and therefore, since ${\color[rgb]{0,0,0}\mathcal{S}}$ is in particular a closed and convex subset of the Hilbert space $\mathcal{H}_{m}$, up to extracting a subsequence there exists $(\widehat{a},\widehat{b})\in{\color[rgb]{0,0,0}\mathcal{S}}$ such that $(a_{k},b_{k})_{k\in\mathbb{N}}$ converges to $(\widehat{a},\widehat{b})$ for the weak topology of $\mathcal{H}_{m}$. Now, thanks to Theorem 2.1 the mapping (18) is continuous for the strong topology of $\mathcal{H}_{m}$. Thus, we infer that it is in particular weakly lower semi-continuous, so that
$\int^{T}_{0}\left\|\frac{\partial\widehat{p}}{\partial t}(t,\cdot)-(\mathcal{L}^{\widehat{a}+\alpha I,\widehat{b}}_{t})^{*}\widehat{p}(t,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}t+\lambda\|(\widehat{a},\widehat{b})\|^{2}_{\mathcal{H}_{m}}\leq$
$\leq\underset{k\to\infty}{\liminf}\ \int^{T}_{0}\left\|\frac{\partial\widehat{p}}{\partial t}(t,\cdot)-(\mathcal{L}^{a_{k}+\alpha I,b_{k}}_{t})^{*}\widehat{p}(t,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}t+\lambda\|(a_{k},b_{k})\|^{2}_{\mathcal{H}_{m}}$
$=\underset{(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m,R_{*}}}}{\min}\quad\int^{T}_{0}\left\|\frac{\partial\widehat{p}}{\partial t}(t,\cdot)-(\mathcal{L}^{a,b}_{t})^{*}\widehat{p}(t,\cdot)\right\|^{2}_{L^{2}}\;\mathrm{d}t+\lambda\|(a-\alpha I,b)\|^{2}_{\mathcal{H}_{m}},$
@@
where the constant $C(a_{*},b_{*})>0$ depends on $a_{*}$ and $b_{*}$ uniquely (in particular, $C(a_{*},b_{*})$ implicitly depends on the hyper-parameters $m$, $p_{0}$, and $\alpha$).
@@
$L_{\lambda}(\widehat{a},\widehat{b})-$ $\lambda\|(a_{*}-\alpha I,b_{*})\|^{2}_{\mathcal{H}_{m}}\leq$
@@
$\leq C(a_{*},b_{*})L(\widehat{p})\leq C(a_{*},b_{*})\left(\log\left(\frac{1}{\delta\varepsilon}\right)^{\frac{1}{2}}\varepsilon\right)^{2},$
@@
We better formalize this metric as follows. For every $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, let $p_{a,b}\in H^{m+1,2(m+1)}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ denote the unique solution to SFPE with coefficients $(\widehat{a},\widehat{b})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. Note that the existence and uniqueness of the regular mapping $p_{a,b}\in H^{m+1,2(m+1)}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ as non-negative solution to SFPE, of unitary mass and with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ is immediate consequence of Theorem 3.2. We define the following metric to test the accuracy of our method:
$E(a,b)\triangleq\|p_{a,b}-p\|^{2}_{L^{2}}=\int^{T}_{0}\|p_{a,b}(t,\cdot)-p(t,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}t,\quad(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}.$